OpenStudy (anonymous):

The basd of a rectangle is 7 cm less than three times its height if the base is 5 cm( what is tge area of the rectangle in square centimeters?

OpenStudy (stefrheart):

First you should put what you know into equations

OpenStudy (nirmalnema):

base=5cm base=3(height)-7 5=3h-7 12=3h h=4 area=5*4 =20cm^2